Define T : R
3 → R
3 by
T(x, y, x) = (−x, x−y,3x+2y+z).
Check whether T satisfies the polynomial (x−1)(x+1)
2
. Find the minimal
polynomial of T.
1
Expert's answer
2019-04-01T11:39:14-0400

Denoting by "I"  the identity operation, we have 

"(T + I) (x, y, z) = (0, x, 3x + 2y + 2z) \, , \\ (T + I)^2 (x, y, z) = (0, 0, 8x + 4y + 4z) \, , \\ (T - I) (x, y, z) = (- 2 x, x - 2 y, 3x + 2y) \, ."


From the last two lines, we then have

"(T - I) (T + I)^2 (x , y , z) = (0 , 0, 0) \, ."

Therefore, "T" satisfies the polynomial "p (x) = ( x - 1 ) ( x + 1 )^2". Considering "v = (1 , 0 , 0) \in \R^3", we have 

"T (v) = (-1 , 1, 3) \, , \quad T^2 (v) = ( 1 , -2 , 2) \, ."

The vectors "v", "T (v)" and "T^2 (v)" are linearly independent; hence, because the space "\R^3" is three-dimensional, the minimal polynomial of "T" has degree three. Since "p (x)" has degree three and is a monic polynomial (the nonzero coefficient of the highest degree is equal to 1), and since "p (T) = 0", it is the minimal polynomial.